OpenStudy (swag):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 72°?

OpenStudy (experimentx):

(n-2)*180/n = 180 - 72, find n, then your sum is (n-2)*180

OpenStudy (ash2326):

Sum of interior and exterior angle is 180 so \[Interior\ angle\ = 180-72=108\] Now we need to find the no of sides of the polygon= n We know that interior angle of n-sides is given by \[\frac{(n-2)\times 180}{n}\] We have \[(n-2)\times 180=108n\] or \[180n-360=108n\] or \[72n=360\] n=5 sides so there are 5 angles each of 108 degrees Sum of the interior angles=108+108+108+108+108=540